Spiller Corp. plans to issue 10%, 9-year, $460,000 par value bonds payable that pay interest semiannually on June 30 and December 31. The bonds are dated December 31, 2016, and are issued on that date. (PV of $1, FV of $1, PVA of $1, and FVA of $1) (Use appropriate factor(s) from the tables provided. Round your "Table value" to 4 decimal places and final answers to nearest whole dollar.)

If the market rate of interest for the bonds is 8% on the date of issue, what will be the total cash proceeds from the bond issue?
Answer #1
Cash proceeds from the issuance of bonds is the present value of cash flow from bond.
Present value of coupon interest $       2,91,164
Present value of Par value $       2,27,056
Present value of cash flows $       5,18,220
So, total cash proceeds from bond is $       5,18,220
Working:
# 1 Semi annual coupon = Par Value * Semi annual coupon rate
= $       4,60,000 * 5%
= $           23,000
# 2 Present value of annuity of 1 = (1-(1+i)^*n)/i Where,
= (1-(1+0.04)^-18)/0.04 i 4%
=             12.6593 n 18
# 3 Present value of 1 = (1+i)^-n
= (1+0.04)^-18
=               0.4936
# 4 Present value of coupon = Coupon payment * present value of annuity of 1
= $           23,000 * 12.6593
= $       2,91,164
# 5 Present value of par value = Par Value * Present value of 1
= $       4,60,000 * 0.4936
= $       2,27,056